OpenStudy (anonymous):

In a math problem I am solving for x and y. I solved for x and got x=2. Now I plug it back into the equation 7x-6y=4. but once I get to -6y = - 10 I don't know what I should write since i cant use a calculator

OpenStudy (oaktree):

Okay. What you do is write it out yourself. Do it like this:\[\frac{ -10 }{ -6 } = \frac{ -1 \times 2 \times 5 }{ -1 \times 2 \times 3 } = \frac{ 5 }{ 3 }\]Do you understand what I did?

OpenStudy (oaktree):

@arielle908?

OpenStudy (anonymous):

what were the -1 for? to make it positive?

OpenStudy (oaktree):

Yeah. If you have a negative on both the top and the bottom mathematicians cancel them out. It's kind of because we're lazy - we don't like to write negatives everywhere if we can help it. :)

OpenStudy (anonymous):

Ohh @OakTree Thank you c:

OpenStudy (oaktree):

No you don't, @Loser66 - she doesn't want to solve the entire problem. And no problem, @arielle908. I was glad to help.

OpenStudy (anonymous):

@Loser66 that was the only part that needed to be answered

OpenStudy (anonymous):

@Loser66 But thank you for looking to help

OpenStudy (oaktree):

Oh, no, that's fine. I was just telling you that the problem was solved. No need to get defensive - it's totally great that you're trying to help out more.

OpenStudy (anonymous):

no worries